MODS1 Commissioning Run 5: UTC 2011 February 22
Observer(s): R. Pogge, J. Antognini, & P. Martini (OSU), M. Wagner (LBTO)
Telescope Operator: D. Gonzalez-Huerta (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: J. Urban (LBTO)
Instrument Support: J. Morris (LBTO)
Heavy ice accumulation on the roof combined with windy conditions (10-18m/s) and cold (-10 to -4°C) all day means that
the snow melt system isn't making much headway. We got the word at 3pm that it was no-go for opening tonight.
End of Commissioning Run 5, and weather is not our friend.
Pity about the ice, the sky above is beautifully clear...
AGw Guide Camera TEC Problem
At the suggestion of Mike Lesser we tested resistance on the MODS guide camera TEC cable at the IUB end. We are reading finite
resistance to the TEC. This means at least that we don't have a break in the circuit, but it does not rule out a short or other
internal TEC problems this test does not address. Updated IssueTrak
#3186 with the test result. We will continue consultation
with Mike and OSU personnel once I get back, and prepare a fix for the March commissioning run.
MODS1 AGw Testing and Telescope Chamber Power Outage Preparation
Shutdown most MODS1 functions in advance of the power outage scheduled for all-day Feb 23:
- MODS1 Red and Blue HEBs and the ICs (quit IC but not power off)
- Calibration Lamps and IMCS laser powered off.
- Reset all linear focus and select mechanisms to their home positions except for the blue and red gratings.
- Wheel mechansims w/o formal reset (filter wheels, grating turrets) set to numerical position 1, and the dichroic to the open position
- Stowed calibration tower and the AGw stage, and made sure the hatch is closed
For the upcoming technical nights where the MODS1 AGw unit will be used, I tested the low-level procedures needed to operate the
stage independent of the MODS1 science instrument. A set of instructions were sent to John Hill.
- 22 Feb 2011